OpenStudy (anonymous):

Let f(x)= 3x+1; g(x)=-2x-5 compute the following

OpenStudy (anonymous):

(f composite g)(10)

OpenStudy (anonymous):

I am a bit confused on how the write this out

OpenStudy (anonymous):

\[f(g(x))=3(-2x-5)+1\] \[f(g(x))=-6x-15+1\] \[f(g(x))=-6x-14\]

OpenStudy (anonymous):

\[f(g(10))=-6(10)-14\]

OpenStudy (anonymous):

g(10)=-25 f(-25)=-74

OpenStudy (anonymous):

how did you get the -25

OpenStudy (anonymous):

g(10), plug in 10 for x in g

OpenStudy (anonymous):

the output of g is then substituted for x in f(x)

OpenStudy (anonymous):

ok that's right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw